# Boundary Attention: Learning to Find Faint Boundaries at Any Resolution

### 01 Jan 2024 v1

@@ 3.3 Visualizing Output
At our network’s output, we expect the shapes of junction boundaries in overlapping regions to agree, so that the variances $\nu_{d}[n],\nu_{f}[n]$ are small at every pixel. Then, the fields of means $\{\bar{d}[n]\},\{\bar{\mathbf{f}}[n]\}$ can be interpreted, respectively, as a global unsigned distance map for the image boundaries and a boundary-aware smoothing of its input channel values. Figure 3 shows an example, where we visualize the zero-set of the global unsigned distance map—we call this the global boundary map—by applying the nonlinearity:
$\bar{b}_{\eta}[n]=\left(1+(\bar{d}[n]/\eta)^{2}\right)^{-1},$ (7)
setting $\eta=0.7$.
The output contains much more information than these fields of means. For any query pixel $n$, we can probe the wedge supports $\{s_{kj}(\cdot;\mathbf{g}[k])\}$ and windowing functions $\{w_{k}(\cdot,\mathbf{p}[k])\}$ that contain it, thereby obtaining a spatial attention map $a_{n}(x)$ that surrounds the query pixel. This is the boundary-aware spatial kernel that turns a neighborhood of input features $\{\mathbf{f}[\cdot]\}$ into the value $\bar{\mathbf{f}}[n]$:
@@
The expression for $a_{n}(x)$ follows from inserting Equation 4 into 6, and its maximum diameter is twice that of $\Omega(x)$. Some examples are shown in the top-right of Figure 3.
Additionally, as shown in the bottom two rows of Figure 3, we can unfold any portion of the output field into the overlapping regions it contains. The figure shows unfolded regions $\{\Omega_{k}(x)\}$ within two windows of the output field. For each unfolded set we visualize: the windowing functions $w_{k}(x;\mathbf{p})$; the regional boundaries $b_{k}(x;\mathbf{g})$ obtained by applying the nonlinearity in Equation 7 to the local distance functions $d_{k}(x;\mathbf{g})$; and the supports $s_{kj}(x;\mathbf{g})$ colored according to the wedge features $\mathbf{f}_{kj}$ that they gather from the input image. In the top example, the neighborhood is homogeneous, so the windowing functions (via $\mathbf{p}[n]$) have large supports, and there are few boundaries (via $\mathbf{g}[n]$). In the bottom example there is fine-scale geometry. The windowing functions are narrow and the primitives agree on a Y-junction.
@@ 4 Network Architecture
We design our network to iteratively refine the fields $\{(\mathbf{g}^{t}[n],\mathbf{p}^{t}[n])\}$. We do so by embedding each field element using a higher dimensional representation, $\boldsymbol{\gamma}^{t}[n]\in\mathbb{R}^{D_{\gamma}}$ and $\boldsymbol{\pi}^{t}[n]\in\mathbb{R}^{D_{\pi}}$ respectively, which can be updated via dot-product attention. In practice, we use $D_{\gamma}=64$ and $D_{\pi}=8$, which provides the network with enough capacity to learn meaningful hidden states. We learn simple linear mappings $\boldsymbol{\gamma}\mapsto\mathbf{g}$ and $\boldsymbol{\pi}\mapsto\mathbf{p}$ which are used for the gather and slice operations. Importantly, all of our network’s elements are invariant to discrete spatial shifts of the image, operating either on individual locations $n$ or on small neighborhoods of locations. See Figure 4 and additional details in the supplement.
@@
Given an input image, the network first applies a neighborhood MLP-mixer, which is a modified variation of MLP-Mixer [32] where we replace the linear operations with convolutions of kernel size $3$. The other change is that we map the input pixels to the hidden state size with a pixel-wise linear mapping rather than taking patches of the input. This block, which we denote the “neighborhood MLP-mixer” transforms the input into an initial hidden state, which is then processed through eight iterations of our boundary attention to further refine the hidden state.
@@
The eight iterations of refinement are broken into two Boundary Attention Blocks with independent weights. To process our input, we first add in a linear projection of the initial hidden state. This is essentially a skip connection that allows our network to retain information from the input pixels at later stages of processing. Next, we copy our hidden state into two identical pieces. We concatenate a dimension 8 learned windowing token to one of the copies and the current estimate of the smoothed global features to the other. We then do neighborhood cross-attention: each pixel in the first copy does two iterations of cross attention with a size 11 patch of the second copy. We add a learned $11\times 1$ 11 positional encoding to the patch, which allows our network to access relative positioning, even if global position cues are absent. We follow each self attention layer with a small MLP.
@@
To transform our output or intermediary hidden state into junction space and render output images, we use a simple linear mapping. We separate the windowing token (the last 8 dimensions) from the hidden state (the first 64 dimensions) and project each through a linear layer. We map the hidden states to 7 numbers that represent $\mathbf{g}=(\boldsymbol{u},\sin(\theta),\cos(\theta),\boldsymbol{\omega})$. These serve as the inputs to our gather and slice operators.
Overall, our network has $2.07\cdot 10^{5}$ learnable parameters, making it orders of magnitude smaller than most learned boundary detectors. For contrast, EDTER [28] has 109 million parameters for its full model.
@@ 4.1 Training
We train our network in two stages: we begin by training the neighborhood MLP-mixer with the first Boundary Attention Block, and then we add the second block and retrain end-to-end. We apply our loss to the two final iterations of our network, so for the first stage of training we apply a loss to iterations 3 and 4, and for end-to-end refinement we apply a loss to iterations 7 and 8. We weigh the final loss three times higher than the second to last loss, which encourages the network to allocate capacity to producing high quality outputs, while ensuring that gradient information is shared across the network.
We train our method using a combination of four global losses applied to global (i.e. averaged) fields, and two patch-wise losses applied to individual patches. The first two losses are supervision losses penalizing mismatches between our network’s predictions and the ground truth feature and boundary maps:
@@
$\alpha[n]=e^{-\beta\cdot(d_{\text{GT}}[n]+\delta)}+C,$ (11)
with $\beta$ and $C$ controlling how much weight to give pixels near boundaries. We set $\beta=0.1$, $\delta=1$, and increase $C$ throughout training to give more weight to non-boundary locations. We also tested a more involved pixel importance mask that gave extra weight to visible vertices and intersections, the details for which can be found in the supplement. Note that using noiseless feature maps for supervision in Equation 9 encourages the windowing functions to be large in smooth regions.
@@
Figure 5: The ODS F-score of our method at different noise levels, compared with the Field of Junctions (FOJ) [33] run using patch sizes $9\times 9$, $17\times 17$, and $25\times 25$, and EDTER [28], with and without preprocessing by BM3D [9]. The bottom insets show example patches at representative PSNR values.
@@ 5 Experiments
We train our model on noisy synthetic data of colorful combinations of triangles and circles. We render $240\times 320$ images containing $15$ to $20$ shapes each, but use $125\times 125$ crops for training. To those crops we add Gaussian and/or Perlin noise [26], and with probability $0.1$ we average over the color channels to produce grayscale inputs. Our dataset contains $10^{5}$ images, $90\%$ of which are used for training, and the rest for testing. For training and optimization details, refer to the supplement.
Performance w.r.t. noise levels.
Figure 5 shows the comparison of our method and baseline approaches, the Field of Junctions [33] and EDTER [28] under different noise levels. We test the Field of Junctions with different patch sizes ($9\times 9$, $17\times 17$, and $25\times 25$) and EDTER with/without preprocessing using optimally-tuned denoising using BM3D [9]. The tuneable parameters for Field of Junctions were chosen to maximize its performance on noisy images with $17\times 17$ patches. Our method outperforms all baselines at lower noise levels and is competitive with the Field of Junctions at higher noise levels, while being orders of magnitude faster than it (see Table 1).
Figure 6: Our network can reliably locate boundaries with sub-pixel precision, due in part to our parametric representation for boundaries. We calculate the ODS F-Score of synthetic images as we increase the pixel matching distance threshold (how close a pixel must be to a ground truth pixel to be counted) from 0 to 1.5 and vary the PSNR of the images from $\infty$ to 8.0. Each line represents the performance at a noise level, with lighter-colored lines corresponding to noisier inputs. X-axis values below 1.0, denoted by the vertical dotted line, represent sub-pixel precision.
Figure 7: One of the advantages of our parametric form for boundaries is that we can trivially upsample our output boundary maps by striding the predicted junction parameters, and scaling the patch size accordingly. This results in clean boundaries even when our output boundary images are upsampled from $125\times 125$ (left) to $500\times 500$ (right).
Figure 8: Top: Linear interpolation in our network’s learned embedding space $\mathbb{R}^{D_{\gamma}}$ from value $\boldsymbol{\gamma}_{a}$ to zero and then to $\boldsymbol{\gamma}_{b}$. Bottom: A geometric interpolation in junction space $\mathbf{g}\in{\cal G}$ that passes through $\mathbf{g}_{0}=(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$. The embedding has learned to be smooth and have an intuitive zero.
Sub-pixel precision.
To measure sub-pixel precision, we render $500\times 500$ images containing pairs of overlapping circles and triangles. This gives us a precise binary boundary map for that resolution. We downsample the images to $125\times 125$, and add varying amounts of Gaussian noise. These noisy, downsampled images serve as the inputs. To evaluate the predicted boundaries, we upsample the outputs back to $500\times 500$ pixels. For EDTER, we use bilinear interpolation along with edge thinning to produce the upsampled boundaries. In our case, a byproduct of our parametric form for boundaries is that an intuitive method for upsampling boundaries naturally arises: by increasing the patch stride to $4$, and rendering the patches as $4$ times as large—68×$\times 68$ 68 pixels—we can create an upsampled version of our image while retaining our method’s accuracy in boundary localization. We visualize this process in Figure 7.
We evaluate the upsampled outputs on the original $500\times 500$ binary ground truth maps and vary the minimum distance threshold of the evaluation metric, which denotes how close a prediction must be to the ground truth to be used. The results are shown in Figure 6. Our F-score remains high across all matching thresholds, even when the inputs contain additive Gaussian noise.
Figure 9: Evolution of boundaries during iterations, in reading order. Early iterations are exploratory and unstructured, while later iterations feature consistent per-patch boundaries, resulting in clean average boundary maps.
Figure 10: Our model generalizes well to real images, and can find accurate boundaries in natural images. Top left: The input image, a low light image from the SIDD [1] dataset. Top right: Our network’s predicted boundaries. Bottom left: The output mean feature map. Bottom right: The predicted distance map.
Linear interpolation in junction space.
Surprisingly, we note that our network learns a spatially smooth manifold of junctions in its hidden state. Figure 8 visualizes some properties of its learned embedding of junction space. We generate equally-spaced samples $\boldsymbol{\gamma}_{i}\in\mathbb{R}^{D_{\boldsymbol{\gamma}}}$ by linearly interpolating from a particular $\boldsymbol{\gamma}_{a}$ to 00 and then to a particular $\boldsymbol{\gamma}_{b}$. We see that the embedding space is smooth, and interestingly, that it learns to associate its zero with nearly-equal angles and a vertex close to the patch center. For visual comparison, we show an analogous interpolation from $\mathbf{g}_{a}$ to $\mathbf{g}_{0}\triangleq(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$ and then to $\mathbf{g}_{b}$ by using a linear geometric interpolation in junction space ${\cal G}$, the expressions for which are in the supplement.
Evolution of the outputs over time.
Figure 9 shows an example of how the distance map $\bar{d}[n]$ evolves during refinement. Specifically, we visualize the result of slicing similar to Equation 5 but with the regional distance functions $d_{k}$ replaced by their nonlinear counterparts $b_{k}$. We see that early iterations are exploratory and unstructured, and that later iterations reach agreement.
Results on real images.
@@
Figures LABEL:fig:real_results and 10 show the qualitative results on real images. Despite being trained on synthetic data, our method can outperform existing state-of-the-art methods at multiple levels of real sensor noise present in ELD [34]. Our method produces crisp and well-defined boundaries despite high levels of noise.
Table 1: Inference time (in seconds) of EDTER [28] (without preprocessing by BM3D [9]), and Field of Junctions (FoJ) [33] run at various patch sizes, compared with our method. We compare both with the original (FoJ) implementation as well as our JAX reimplementation (JAX-FoJ). Unlike FoJ [33], our method does not require striding for fitting moderately-sized images in memory (runtimes reported with $*{}^{*}$ did require striding). The average run times were calculated on an Nvidia A100 GPU.
$125\times 125$ $320\times 320$
EDTER [28] $0.130$ $0.130$
FoJ [33]  $9\!\!\times\!\!9$ $36.8$ $206$
 $17\!\!\times\!\!17$ $88.0$ $76.0^{*}$ (stride-$3$)
 $25\!\!\times\!\!25$ $162$ $57.6^{*}$ (stride-$5$)
@@
Ours $0.0823$ $0.678$
Inference time.
Table 1 compares the running time for different methods at two resolutions. We denote several runs for Field of Junctions where we had to increase the patch stride (optimize junctions at strided intervals) to avoid running out of memory. Notably, we are orders of magnitude faster than Field of Junctions, despite achieving similar performance. Additionally, because our network can handle inputs of any size, the run time depends on the input shape, in contrast with EDTER that has a maximal input size of $320\times 320$ inputs and constant run time.

### 18 Mar 2024 v2

@@ 3.3 Visualizing Output
At our network’s output, we are provided with a field of tuples representing the junction and window parameters $\{(\mathbf{g}[k],\mathbf{p}[k]\}$ for all stride-1 patches of the input image. Though these outputs are defined continuously, we can visualize them in Figure 3 by rendering the windowing functions $w_{k}(x;\mathbf{p})$ (second column), and the binary junction wedge supports $s_{kj}(x;\mathbf{g})$ colored purple, orange, and blue (third column), along with the input image patches (first column). To the left, we show the input image and the regions from which the patches were extracted from.
Additionally, we can use the learned junction parameters to render unsigned distance patches $d_{k}(x;\mathbf{g})$ (fourth column), boundary patches $b_{k}(x;\mathbf{g})$ (fifth column), and the binary junction wedges $s_{kj}(x;\mathbf{g})$ colored with their respective gathered colors (sixth column). All of the patches are defined continuously, so we can render them at any resolution.
We expect the shapes of junction boundaries in overlapping regions to agree, so that the variances $\nu_{d}[n],\nu_{f}[n]$ are small at every pixel. Then, the fields of means $\{\bar{d}[n]\},\{\bar{\mathbf{f}}[n]\}$ can be interpreted, respectively, as a global unsigned distance map for the image boundaries (bottom row, fourth column) and a boundary-aware smoothing (bottom row, last column) of its input channel values. Figure 3 shows an example (bottom row, fifth column), where we visualize the zero-set of the global unsigned distance map—we call this the global boundaries—by gathering boundary patches $b_{k}$ defined as:
$b_{k}(x;\mathbf{g})=\left(1+(d_{k}(x;\mathbf{g})/\eta)^{2}\right)^{-1},$ (7)
setting $\eta=0.3$.
The output contains much more information than these fields of means. For any query pixel $n$, we can probe the wedge supports $\{s_{kj}(\cdot;\mathbf{g}[k])\}$ and windowing functions $\{w_{k}(\cdot,\mathbf{p}[k])\}$ that contain it, thereby obtaining a spatial attention map $a_{k}(x)$ that surrounds the query pixel. This is the boundary-aware spatial kernel that turns a neighborhood of input features $\{\mathbf{f}[\cdot]\}$ into the value $\bar{\mathbf{f}}[n]$:
@@
The expression for $a_{n}(x)$ follows from inserting Equation 4 into 6, and its maximum diameter is twice that of $\Omega(x)$. The spatial affinity maps for two probed points are shown in the leftmost image of the bottom row of Figure 3. Like slicing, probing the affinity map around a pixel is a form of global accumulation that takes values defined per-patch and synthesizes them into a global representation.
@@ 4 Network Architecture
We seek a differential architecture that can effectively initialize and then refine the fields $\{(\mathbf{g}^{t}[k],\mathbf{p}^{t}[k])\}$ using a manageable number of iterations. This bears similarities to the edge localization and edge-linking steps of early edge detectors like Canny’s, but with a more sophisticated local boundary model, and with mechanisms for spatial consistency that are learned instead of designed. It is also analogous to the original field of junctions algorithm [25], which uses coordinate descent for initialization and iterations of gradient descent for refinement. We want to replace both steps with something that is differentiable, faster, and able to scale to larger images.
After considering a variety of alternatives, we settled on a particular family of architectures based on dot-product self-attention; and within this family, we performed an extensive set of ablations to determine the best performing model. We describe our final model here, and we provide the ablation details in Section 6. Importantly, all of our model’s components are invariant to discrete spatial shifts of the image, operating either on individual locations $k$, or on small neighborhoods of locations with spatially-shared parameters. This means that our model can be trained on small images and then deployed on much larger ones. Also, our model has only $207$k learnable parameters, making it orders of magnitude smaller than most deep semantic boundary detectors. As a point of reference, the full EDTER model has 109M parameters [19].
Our model is depicted in Figure 4. It implicitly uses higher-dimensional embeddings of the field elements, $\boldsymbol{\gamma}^{t}[k]\in\mathbb{R}^{D_{\gamma}}$ and $\boldsymbol{\pi}^{t}[k]\in\mathbb{R}^{D_{\pi}}$, which can be decoded at any iteration using the learned linear mappings $\boldsymbol{\gamma}\mapsto\mathbf{g}$ and $\boldsymbol{\pi}\mapsto\mathbf{p}$. Our final model uses $D_{\gamma}=64$ and $D_{\pi}=8$ which, as shown in the supplement, provides enough capacity to learn a smooth latent representation of junction space. Use separate embeddings for the junction and windowing fields provides a disentangled representation of both.
@@
Given an input image, the network first applies a “neighborhood MLP-mixer”, which is a modified MLP-Mixer [24] that replaces the global spatial operations with convolutions of kernel size $3$. The other change is that we map the input pixels to the hidden state size with a pixel-wise linear mapping rather than taking patches of the input. This block transforms the input image into a pixel-resolution map with ${D_{\gamma}}$ channels. We denote this by $\gamma_{0}[n]$ and refer to it as the initial “hidden state”. This hidden state is then refined by a sequence of eight boundary attention iterations, which we describe next. (See the supplement for a visualization of the decoded hidden states as they evolve.)
@@
The eight iterations of refinement are broken into two boundary attention blocks, each with learned weights. To process our input, we first add in a linear projection of the initial hidden state. This is essentially a skip connection that allows our network to retain information from the initial hidden state estimate at later stages of processing. Next, we copy our hidden state into two identical pieces. We concatenate a dimension 8 learned windowing embedding to one of the copies and the input image plus the current estimate of the smoothed global features to the other. We then do neighborhood cross-attention: each pixel in the first copy does two iterations of cross attention with a size 11 patch of the second copy. We add a learned $11\times 1$ 11 positional encoding to the patch, which allows our network to access relative positioning, even though global position cues are absent. We follow each self attention layer with a small MLP.
@@
To transform our output or intermediary hidden state into junction space and render output images, we use a simple linear mapping. We separate the windowing embedding (the last 8 dimensions) from the junction embedding (the first 64 dimensions) and project each through a linear layer. We map the hidden state to 7 numbers that represent $\mathbf{g}=(\boldsymbol{u},\sin(\theta),\cos(\theta),\boldsymbol{\omega})$. These serve as the inputs to our gather and slice operators.
To extract $\mathbf{p}$ from the windowing embedding, we linearly project the windowing embedding to a length 3 vector, which we use as coefficients in a weighted sum of three square pillbox functions with widths $3$, $9$, and $17$. This implementation of the windowing function ensures spatial overlap between neighboring patches by limiting the minimum patch size to $3$.
@@ 4.1 Training
We train our network in two stages: we begin by training the neighborhood MLP-mixer with the first Boundary Attention Block, and then we add the second block and retrain end-to-end. Because junction space is non-convex and difficult to optimize over, this two step training procedure helps the network converge.
We apply our loss to the two final iterations of our network, so for the first stage of training we apply a loss to iterations 3 and 4, and for end-to-end refinement we apply a loss to iterations 7 and 8. We weigh the final loss three times higher than the second to last loss, which encourages the network to allocate capacity to producing high quality outputs, while still providing supervision to intermediate junction estimates.
We train our method using a combination of four global losses applied to global (i.e. sliced) fields, and two patch-wise losses applied to individual patches. The first two losses are supervision losses penalizing mismatches between our network’s predictions and the ground truth feature and boundary maps:
@@
$\alpha[n]=e^{-\beta\cdot(d_{\text{GT}}[n]+\delta)}+(1-C),$ (11)
with $\beta$ and $C$ controlling how much weight to give pixels near boundaries. We set $\beta=0.1$, $\delta=1$, and anneal $C$ to give more weight to non-boundary locations later in training. We also tested a more involved pixel importance mask that gave extra weight to visible vertices and intersections, the details for which can be found in the supplement. Note that using noiseless feature maps for supervision in Equation 9 encourages the windowing functions to be large in smooth regions.
@@
Figure 5: Left: ODS F-score for our method and multiple baselines at different noise levels computed on noisy synthetic data. The bottom inset show example patches at representative PSNR values. Our method outperforms all baselines at low noise and is better or competitive with other techniques at high noise. Right: Comparing the F-score for different techniques with their runtime. Our method has the best average F-score while also being much faster than the second best method Field of Junctions.
Figure 6: Left: Qualitative comparison of our model versus other methods for noisy crops from the ELD dataset [26]. Our method provides more detail than other techniques at low noise and is more robust to high levels of noise. Right: Repeatability of the estimated boundaries of crops from the ELD dataset over increasing noise. Over multiple levels of real sensor noise, our method is the most consistent in predicting edges.
@@ 5 Experiments
We train our model on noisy synthetic data of colorful combinations of triangles and circles. We render $240\times 320$ images containing $15$ to $20$ shapes each, but use $125\times 125$ crops for training. To those crops we add Gaussian and/or Perlin noise [17], and with probability $0.1$ we average over the color channels to produce grayscale inputs. Our dataset contains $10^{5}$ images, $90\%$ of which are used for training, and the rest for validation. For a visualization of training data and training and optimization details, refer to the supplement.
Baselines.
Since we focus on low-level edge detection, we primarily compare against other techniques with a similar focus and that also do not train on large semantic datasets: Canny [3], gPb [14], Structured Edges (SE) [7], and Field of Junctions (FoJ) [25]. However, since most recent work tends to focus on semantic edge detection, in the interest of being comprehensive, we also provide a comparison to several modern works: UAED [28], EDTER [19], and HED [27]. We note that this is not a clean comparison due to the difference in tasks – predicting edges with low-level cues vs predicting edges with semantic meaning.
Quantitative results and inference times across different noise levels.
For quantiative evaluation, our focus on low-level cues means that we cannot use semantic edge detection benchmarks like BSDS50 [2] since our method and other low-level baselines will both predict edges that are not semantically meaningful and fail to predict edges that are semantically meaningful but lack contrast. We instead rely on synthetic data where the ground truth edges can be determined with perfect precision and any level of noise can be added. In this experiment, we use geometric objects and per-pixel additive Gaussian noise with a random variance.
Figure S3 compares the performance and inference time of our method and baselines under different noise levels. The tuneable parameters for Field of Junctions were chosen to maximize its performance on noisy images with $17\times 17$ patches. Notably our method’s adaptive windowing function gives it at edge compared to the Field of Junctions at low noise, enabling it to capture finer details, with only slightly worse performance under extreme noise conditions. Our method is also orders of magnitude faster than FoJ as shown in the plot to the right.
Qualitative results on real images.
@@
As shown in Figure 6, despite being trained on synthetic data, our method can outperform existing state-of-the-art methods at multiple levels of real sensor noise present in ELD [26]. Our method produces crisp and well-defined boundaries despite high levels of noise. Figure 7 shows that our method makes reasonable boundary estimates for even complex in-the-wild images. Our supplement contains additional examples.
Figure 7: Our model can find compelling boundaries in natural in-the-wild images, despite training on simple synthetic data. Left image from the SIDD [1] dataset; right image taken with a phone camera.
Repeatability on real images.
We quantify each model’s noise resilience on real low-light images by measuring the repeatability of its boundaries over a collection of scenes from the ELD dataset. For each scene, we run the model on the lowest-noise image and then measure, via ODS F-score, how much its predictions change with increasing noise level. Figure 6 shows the averages of these scores across the collection. Our model provides more consistent results for increasingly noisy images than other methods.
6 Ablations
[…]
table:ablation1 shows what happens when the initial per-patch Neighborhood MLP-mixer [NM] is used alone, versus combining it with one boundary attention block [BA-1] or two such blocks [BA-2]. Tying the attention weights across iterations provides a slight advantage at higher noise levels, with only a slight penalty in accuracy at lower noise. Our final model is boxed. It provides the best overall performance, accepting slightly lower accuracy at low noise in exchange for better accuracy at higher noise levels.
@@ 6 Ablations
N/A $0.355$ $0.189$ $0.179$
$0.837$ $0.582$ $0.293$
0.874 $0.658$ $0.327$
$0.872$ 0.673 0.348
Table 1: Ablations: impact of component combinations and weight-tying on F-score (higher is better). Our reported model is boxed.
[…]
Iterations 2 $0.857$ $0.611$ $0.307$
3 $0.867$ $0.647$ $0.329$
4* 0.872 0.673 0.348
5 $0.871$ $0.663$ $0.337$
Neighborhood $3\times 3$ $0.835$ $0.531$ $0.240$
$7\times 7$ $0.869$ $0.637$ $0.313$
$9\times 9$ $0.871$ $0.658$ $0.325$
$11\times 1$ 11* 0.872 0.673 0.348
$13\times 13$ $0.871$ $0.655$ $0.324$
MLP Input Constant $0.870$ $0.650$ $0.317$
Input features $0.868$ $0.657$ $0.316$
Avg. features* 0.872 0.673 0.348
Loss Uniform $0.844$ $0.629$ $0.313$
Weighted* 0.872 0.673 0.348
Windowing Fixed - $3\times 3$ $0.834$ $0.546$ $0.205$
Fixed - $9\times 9$ 0.872 $0.662$ $0.328$
Fixed - $17\times 17$ $0.859$ $0.662$ $0.328$
Inferred* 0.872 0.673 0.348
Table 2: Ablations: variations of our reported model. We compare the number of iterations of boundary attention, the attention neighborhood size, the features concatenated to the hidden state as input to the MLP, the impact of a spatially weighted loss, and a fixed versus learned windowing function. Asterisks indicate our reported choices.